The angle of acceptance of an optical fibre is 30 degree when kept in air. find the angle of acceptance when it is in a medium of refractive index 1.33
1
Expert's answer
2020-05-25T10:47:32-0400

Acceptance angle= "\sin^{-1}(\frac{NA}{n_o})=\sin^{-1}(\frac{n_1^2-n_2^2}{n_o})"

in air "n_o=1"

"30^o=\sin^{-1}(NA)"

NA=0.5


acceptance angle when refractive index of medium is 1.33

new acceptance angle = "\sin^{-1}(\frac{0.5}{1.33})=\sin^{-1}0.37" ="21.7^o"